OpenStudy (cutiekawdfi):

will fan and medal :) Find the foci of each hyperbola 1. y^2/400 - x^2/169 = 1

OpenStudy (cutiekawdfi):

@Zarkon

OpenStudy (cutiekawdfi):

@Tommynaiter

OpenStudy (anonymous):

y^2/400- x^2/169= 400=a^2 169=b^2 a^2+b^2=c^2 so 400+169=c^2 \[7\sqrt{569}\] =c answer (0,\[7\sqrt{569}\] and (0,-\[\sqrt{569}\])